# Population of tetraneutron continuum in reactions of 8He on deuterium

### v5

@@ 3 On 4n population in different reactions
The low-energy population in [6] and in the processes considered in this work all have the 8He nuclide as a starting point. At first glance, this is the only similarity, and the reaction mechanisms should be too different. Let’s demonstrate that the differences are actually not so striking. Table 1 shows the momenta and velocities in the recoil-product channel for the reaction of Ref. [6] and the reactions considered in this work. Of course, in the reaction of [6] the “recoil” particles are leaving the FSI region faster, than in the case of our Li) reaction, but in the timescale this “faster” is only a factor of 2.
@@
| 7HH+ | 3H |
The 2H(8He,3He)7HH+ process formally has the decay process as the second step. In fact, from Table 1 it is clear that for the excited states of 7H, which we consider, the velocity of the 3H recoil leaving 4n in the “decay process” is actually comparable with the velocity, at which the 3He recoil had left the FSI region in the first [the 2H(8He,3He)7H reaction] step of this process. This kinematical region is actually a transition region from 2H(8He,3He)7H reaction to 2H(8He,n, where highly excited states of 6Li (those located above the 3He-3H threshold) are populated. We find in Fig. 4 (c) that the most expressed low-energy 4n hump not “contaminated” with higher-energy phase-volume-like contributions is obtained for the recoil, which is the lowest 3He-3H resonant state.
So, we see that the low-energy 4n continuum is confidently populated in the 2H(8He,n reactions with excitation energies MeV, see Fig. 2 (b), and MeV, see Fig. 4 (c). It is quite natural to expect that such a population should also take place at the intermediate energies, which correspond to + continuum states of 6Li. These would be reactions of quasifree scattering 2H(8He,)4n, which have never considered seriously because of extremely “soft” 2H target. So far, the possibility to extract this information from our existing data is questionable.
Another characteristic, which is expectedly very important for the direct reactions, is momentum transfer . For our reactions the values are easily reconstructed, see Fig. 5. The momentum transfers MeV are realized, what is found important for data interpretation, see Fig. 7 and related text. Unfortunately, we do not find information on the momentum transfers in Ref. [6], though this information is obviously accessible in these data.
4 Theoretical model
@@
The simpliest reaction model which can be considered for 4n production is “sudden removal” of core from 8He [19, 23]. This procedure defines the source function in (2) as the Fourier transform of the overlap integral of the -cluster WF and the 8He WF over the radius-vector between the removed -cluster and the 4n center-of-mass:
The “strength function” for population of the 4n continuum should be proportional to the outgoing flux of particles, expressed via the WFs with pure outgoing wave boundary conditions Eq. (5)
@@ 5 Properties of the 8He
In the calculations we adopted the local charge independent Argonne SSC AV14 potential [38]. For evaluations we also used simple -wave singlet potential Brown-Jackson (BJ) from [39]: , MeV, fm. For three nucleons with separations we use the three-nucleon potential [40]
@@
6 The 8He WF as a source for 4n population
The 8He WF correlation density vs. (informative for the 4n studies) is shown in Fig. 6 (a). It looks overall quite featureless. In contrast, the projection of the 8He WF on the 4n configuration with , where strongest 4n FSI effect is expected, exhibits strong spatial correlations, see Fig. 6 (b). The correlation density vs. can be easier perseived basing on a simple relation between the rms radius of individual neutron in the 4n c.m. frame and the 4n rms hyperradius
@@ 6 The 8He WF as a source for 4n population
It should be understood that the strongly correlated source shown in Fig. 6 (b) is related to about of the 8He g.s. WF. The other components of these WF projected on 4n system populate “excited” configurations of 4n, where we do not expect any strong final-state effect.
Figure 7 (a) shows the evolution of 4n strength functions for selected MeV: no FSI reduced FSI full 4n FSI. The strength function peak for full 4n FSI is at about 14 MeV in a good agreement with analogous calculation in [19]. To form extreme low-energy peak in the strength function in the region MeV as indicated by data, a further qualitative assumption of the peripheral character of reaction is needed. To evaluate it, we cut out small hyperradii of the original source function by a profile function of the form
@@
The systematics of the strength functions for peripheral sources for different momentum transfers is illustrated in Fig. 7 (b). One may see that even the “peripheral assumption” is not sufficient to form a strong and distinct low-energy peak in the 4n spectrum: the desired picture is realized only for certain intermediate range of transferred momenta MeV. This range well corresponds to situation of our experiment, see Fig. 5 and the related discussion in Section 3.
@@
Qualitative comparison of theory and available data is provided in Fig. 8. Here we need to recall that strongly correlated picture of Fig. 7 is connected with of the 8He WF projecting on the state of 4n, while of this WF populate excited configurations of the 4n. The latter WF components are evaluated to provide strength function with maximum at MeV, see thick gray curve in Fig. 8 (d). As we have mentioned in Section 3, the information on transfered momenta is not available in Ref. [6]. This means that, so far, we can not find out even on the qulitative level, whether the experimental results of this work should be in principle, the same as in [6] (and they just seem somewhat different only because of experimental resolutions and statistics), or there should be some significant physical difference between them due to difference in the reaction mechanisms.
It is also important to note that the low-energy peak connected with extreme peripheral source has very different profile compared to the case of the real resonant behavior connected with normal-size source. The “resonant” (red dashed) curve in Fig. 8 (d) was generated by additional artificial binding introduced in Hamiltonian by phenomenological 4-body potential (11) with parameters fm and fm. The difference between “peripheral” black and “resonant” red curves in Fig. 8 (d) is too small to be seriously discussed for the available quality of the data. However, in principle, we have guidelines from theory how to distinguish experimentally the “peripheral” (caused by reaction mechanism) and “true resonant” tetraneutron continuum responces.
7 Theoretical discussion
@@
So, in principle, there exist a class of source functions which can produce extreme low-energy peak in the 4n spectrum for quite compact ( fm) source. The only problem is that the “supersoft” source Eq. (12) has short-distance asymptotic . If we get back to the definition Eq. (6), this implies that 8He overlap with 4-body hypersphericals is singular as at zero, which is highly unlikely for any realistic 8He structure. The 4n sources deduced in our work has asymptotic at zero for . This is in full agreement with analytical results of [19]. Consequently, such sources are much “harder” in the momentum space, and additional peripheral assumption is needed to get sizable low-energy effect in 4n spectrum.
The origin of the “supersoft” source in [23] is probably due to the fact that 4n c.m. is assumed to coincide with 8He c.m. in Ref. [23]. This is an implicit assumption of infinitely heavy -core in 8He and, of course, quite a poor approximation considering the actual mass ratio of and 4n. This means that within the cluster model the charge radius of 8He coincides with the charge radius of the -particle, which contradicts to the data [43]. This assumption also leads to total absence of such an important characteristic as in the formalism of [23]. In contrast, the “geometric” effects of 4n motion relative to in 8He WF are found to be very pronounced in our calculations, see Fig. 6. Also the rms size of the source fm, see e.g. Fig. 7 (a), presume significantly smaller fm distances in 4n source, than the distances in the initial 8He WF, see Table 2.
@@
So, we insist that peripheral assumption is necessary to describe the extreme low-energy peak in the spectrum of 4n. This point is also qualitatively supported by extreme small cross sections for the observed peculiarities in the low-energy 4n spectra. Typical cross section values for analogous direct reactions are some millibarns and even tens of millibarns, while we and Ref. [6] observe some microbarns. Such a cross section suppression looks quite natural if only the extreme periphery of the original 4n configuration is actually participating in the reactions of interest without destroying rescattering effects.
8 Conclusion
@@
Evidence for a hump in the continuum at and MeV was observed in the two reactions: 2H(8He,6Li) and 2H(8He,3He)7HH+, respectively. The peak energies in our work and MeV in Ref. [6] are also consistent within experimental resolutions. We demonstrate that the reaction mechanism in both cases is actually best interpreted as transfer reaction to the ground and to the MeV states of 6Li. The reaction mechanism also can be found reasonably consistent with that in the 1H(8He, 4He) reaction [6], so it could be the same type of phenomenon. The statistics of 6 events for the first our reaction is quite low, but events for the second one is comparable to statistics connected with the low-energy structure observed in [6].
@@
Paper [23] is explaining the low-energy tetraneutron peak in Ref. [6] as solely an initial state (8He structure) effect in the presence of 4n FSI. This is in a strong contradiction to the results of [19] and the results of this work. As compared to [19] the calculations of this work (i) are performed with realistic and interactions, and (ii) the obtained realistic 8He source demonstrates the complicated correlated behavior that is very important for the 4n population. So far, the strong spatial correlations (Pauli focusing) for the 8He g.s. WF has been several times discussed in the qualitative models [45, 46, 31], but in our work they are for the first time explicitly demonstrated in the realistic calculations of 8He g.s. structure.
@@
The two aspects mentioned above are not sufficient to fully explain the presence of a low-lying peak in the tetraneutron spectrum. To do this, it is also necessary to consider extremely peripheral reaction mechanism, in which the -core of 8He interacts with the target, but four neutrons remain “spectators” at distances, far exceeding the typical sizes of neutron orbitals in 8He. Using the appropriate profile function of the type Eq. (10), simulating such a geometry, it is possible to explain the presence of a low-lying peak in the tetraneutron spectrum. Since the appearance of the low-energy peak is not related to the tetraneutron per se, but also to the ISS (the source used) and the reaction mechanism, it should be expected that such a structure could be detected at somewhat different energies in different reactions. A possible reason of disagreement with the calculations [23] may be the incorrect 4n c.m. treatment in the 8He WF used in [23].
@@
It is clear that the further studies of 8He induced reactions on the deuteron target are needed to bring to full reliability the evidences obtained in this work. However, the processes of the major interest are all in the cross sections range of microbarns. Therefore, such studies make sense only if the statistical limitations of the existing data are overcome by about an order of the magnitude, which is a real challenge for available experimental techniques and facilities.

### v8

@@ III On tetraneutron population in different reactions
The low-energy population of 4n in Ref. [6] and in the processes considered in this work all have the 8He nuclide as a starting point. At first glance, this is the only similarity, and the reaction mechanisms should be too different. Let’s demonstrate that the differences are actually not so striking.
Table 1 shows the momenta and velocities in the recoil-product channel for the reaction of Ref. [6] and the reactions considered in this work. Of course, in the reaction of [6] the “recoil” particles are leaving the FSI region faster, than in the case of our Li) reaction, but in the timescale this “faster” is only a factor of 2.
@@
| 7HH+4n | 3H |
The 2H(8He,3He)7HH+4n process formally has the 7HH+4n decay process as its second step. In fact, from Table 1 it is clear that for the relatively highly-excited states of 7H, which we consider, the velocity of the 3H recoil leaving 4n in the “decay process” is actually comparable with the velocity, at which the 3He recoil had left the FSI region at the first [the 2H(8He,3He)7H reaction] step of this process. This kinematical region is actually a transition region from 2H(8He,3He)7H reaction to 2H(8He,n reaction, where highly excited states of 6Li (those located above the 3He-3H threshold) are populated. We find in Fig. 4 (c) that the most pronounced low-energy 4n hump, not “contaminated” by higher-energy phase-volume-like contributions, is obtained for the recoil, which is the lowest-energy resonant state in the 3He-3H continuum.
So, we see that the low-energy 4n continuum is confidently populated in the 2H(8He,n reactions with excitation energies MeV, see Fig. 2 (b), and MeV, see Fig. 4 (c). It is quite natural to expect that such a population should also take place at the intermediate energies, which correspond to + continuum states of 6Li. In this case we deal with reaction of quasifree scattering 2H(8He,)4n, which has never been considered seriously because of extremely “soft” nature of the 2H target. Unfortunately we were not able to extract this information from our existing data due to technical limitations.
Another characteristic, which is expectedly very important for the direct reactions, is momentum transfer to the 4n system, see, e.g. the Eq. (9). For our reactions the reconstructed values are shown in Fig. 5. Quite different momentum transfers and MeV are realized for the low-energy 4n population in the 2H(8He,6Li) and 2H(8He,3He)7HH+4n reactions, respectively. This is found important for data interpretation, see Fig. 7 and related text. Unfortunately, we do not find information on the momentum transfers in Ref. [6], though this information is obviously accessible in these data. Only the restriction MeV can be roughly estimated from Fig. 2 of Ref. [6], while the most probable value seem to be around MeV.
IV Theoretical model
@@
The simplest reaction model which can be considered for 4n production is “sudden removal” of core from 8He [19, 23]. This procedure defines the source function in (4) as the Fourier transform of the overlap integral of the -cluster WF and the 8He WF over the radius-vector between the removed -cluster and the 4n center-of-mass:
Within the sudden removal approximation the momentum transferred to the 4n system is equal to the momentum of the removed -cluster. The “strength function” for population of the 4n continuum should be proportional to the outgoing flux of particles, expressed via the WFs with pure outgoing wave boundary conditions Eq. (7)
@@ V Properties of the 8He
In the calculations we adopted the local charge-independent Argonne SSC AV14 potential [42]. For evaluations we also used simple -wave singlet potential Brown-Jackson (BJ) from [43]: , MeV, fm. For three nucleons with separations we use the three-nucleon potential [44]
@@
VI The 8He WF as a source for tetraneutron population
The 8He WF correlation density vs. (collective radius of four neutrons, informative for the 4n studies) is shown in Fig. 6 (a). It looks overall quite featureless. In contrast, the projection of the 8He WF on the 4n configuration with , where strongest 4n FSI effect is expected, exhibits strong spatial correlations, see Fig. 6 (b). The correlation density vs. can be easier perceived basing on a simple relation between the rams radius of individual neutron in the 4n c.m. frame and the 4n rams hyperradius
@@ VI The 8He WF as a source for tetraneutron population
It should be understood that the strongly correlated source shown in Fig. 6 (b) is related to about of the 8He g.s. WF. The other components of these WF projected on 4n system populate “excited” configurations of 4n, where we do not expect any strong FSI effect.
Figure 7 (a) shows the evolution of 4n strength functions for one selected MeV: no FSI reduced FSI full 4n FSI. The strength function peak for full 4n FSI is at about 14 MeV in a good agreement with analogous calculation in [19]. To form extreme low-energy peak in the strength function in the region MeV as indicated by data, a further qualitative assumption of the peripheral character of reaction is needed. To evaluate it, we cut out small hyperradii of the original source function by a profile function of the form
@@
The systematics of the strength functions for peripheral sources for different momentum transfers is illustrated in Fig. 7 (b). One may see that even the “peripheral assumption” is not sufficient to form a strong and distinct low-energy peak in the 4n spectrum: the desired picture is realized only for certain intermediate range of transferred momenta MeV. This range well corresponds to situation of our experiment, see Fig. 5 and the related discussion in Section III.
@@
Qualitative comparison of theory and available data is provided in Fig. 8. Here we need to recall that strongly correlated picture of Fig. 7 is connected with of the 8He WF projecting on the state of 4n, while of this WF populate excited configurations of the 4n. The latter WF components are evaluated to provide strength function with maximum at MeV, see thick gray curve in Fig. 8 (d). As we have mentioned in Section III, the information on transferred momenta is not available in Ref. [6], only the rough upper limit estimate MeV. This means that, so far, we can not find out even on the qualitative level, whether the experimental results of this work should be in principle, the same as in [6] (and they just seem somewhat different only because of experimental resolutions and statistics), or there should be some significant physical difference between them due to difference in the reaction mechanisms connected with value.
It is also important to note that the low-energy peak connected with extreme peripheral source has very different profile compared to the case of the real resonant behavior connected with normal-size source. The “resonant” (blue dashed) curve in Fig. 8 (d) was generated by additional artificial binding introduced in Hamiltonian by phenomenological 4-body potential (13) with parameters fm and fm. The difference between “peripheral” black solid and “resonant” blue dashed curves in Fig. 8 (d) is too small to be seriously discussed for the available quality of the data. However, in principle, we have guidelines from theory how to distinguish experimentally the “peripheral” (caused by reaction mechanism) and “true resonant” tetraneutron continuum responses.
VII Theoretical discussion
@@
So, in principle, there exist a class of source functions, which can produce extreme low-energy peak in the 4n spectrum for quite a compact ( fm) source. The only problem is that the “supersoft” source Eq. (14) has short-distance asymptotic . If we get back to the definition Eq. (8), this implies that 8He overlap with 4-body hypersphericals is singular as at zero, which is highly unlikely for any realistic 8He structure. The 4n sources deduced in our work has asymptotic at zero for ; they can be reasonably approximated by expression like (14), but only for . This is in full agreement with analytically deduced source functions of Ref. [19]. The source with is evidently much “harder” in the momentum space than the source of the same radius with . Therefore, an additional peripheral assumption is needed to get sizable low-energy effect in 4n spectrum with such a source.
So, we guess that the low-energy enhancement of the 4n spectrum in Ref. [23] is connected with “supersoft” source. The origin of the “supersoft” source in this work is probably due to the fact that 4n c.m. is assumed to coincide with 8He c.m. in Ref. [23] leading to several important inconsistencies:
(i) This is an implicit assumption of infinitely heavy -core in 8He and, of course, quite a poor approximation considering the actual mass ratio of and 4n.
(ii) This also means that within the cluster model the charge radius of 8He coincides with the charge radius of the -particle, which contradicts to the experimental data [47].
(iii) The “infinitely heavy core” assumption also leads to total absence of such an important characteristic as in the formalism of [23]. In contrast, the “geometric” effects of 4n motion relative to in 8He WF are found to be very pronounced in our calculations, see Fig. 6.
(iv) The rms size of the source fm, see e.g. Fig. 7 (a), presume significantly smaller fm distances in the 4n source, than the fm distances in the initial 8He WF, see Table 2. The “infinitely heavy core” assumption leads to equal distances in the 4n source and in the initial 8He WF.
@@
So, we insist that peripheral assumption is necessary to describe the extreme low-energy peak in the spectrum of 4n. This point is also qualitatively supported by extreme small cross sections for the observed peculiarities in the low-energy 4n spectra. Typical cross section values for analogous direct reactions are some millibarns and even tens of millibarns, while we and Refs. [5, 6] observe some microbarns and even nanobarns. Such a cross section suppression looks quite natural if only the extreme periphery of the original 4n configuration of 8He is actually participating in the reactions of interest without destructive rescattering effects.
VIII Conclusion
@@
Evidence for a hump in the continuum at and MeV was observed in the reactions 2H(8He,6Li)4n and 2H(8He,3He)7HH+4n, respectively. The 4n peak energies in our work and the “resonance-like structure” at MeV observed in Ref. [6] are consistent within experimental resolutions. We demonstrate that the reaction mechanism in both reported here reactions is actually best interpreted as transfer reaction to the ground state and to the MeV excited state of 6Li, respectively. The reaction mechanism also can be found reasonably consistent with that in the 1H(8He, 4He)4n reaction [6], so it could be the same type of phenomenon. The statistics of 6 events for the first our reaction is quite low, but events for the second one is comparable to statistics connected with the low-energy 4n resonance-like structure observed in [6].
@@
Paper [23] is explaining the low-energy tetraneutron peak in Ref. [6] as solely an initial state (8He structure) effect in the presence of 4n FSI. This is in a strong contradiction to the results of [19] and the theoretical results of this work. As compared to [19] the calculations of this work (i) are performed with realistic and interactions, and (ii) the obtained realistic 8He source demonstrates the complicated correlated behavior that is very important for the 4n population. So far, the strong spatial correlations (Pauli focusing) for the 8He g.s. WF has been several times discussed in the qualitative models [49, 50, 35], but in our work they are for the first time explicitly demonstrated in the realistic calculations of 8He g.s. structure.
@@
The two aspects mentioned above are not sufficient to fully explain the presence of a low-lying peak in the tetraneutron spectrum. To do this, it is also necessary to consider extremely peripheral reaction mechanism, in which the -core of 8He interacts with the target, but four neutrons remain “spectators” at distances, far exceeding the typical sizes of neutron orbitals in 8He. Using the appropriate profile function of the type Eq. (12), simulating such a geometry, it is possible to explain the presence of a low-lying peak in the tetraneutron spectrum. Since the appearance of the low-energy peak is not related to the tetraneutron per se, but also to the ISS (the source used) and the reaction mechanism, it should be expected that such a structure could be detected at somewhat different energies in different reactions. It is demonstrated that a likely reason of disagreement with the calculations [23] may be the incorrect 4n c.m. treatment in the 8He WF used in [23].
@@
It is clear that the further studies of 8He induced reactions on the deuteron target are needed to bring to full reliability the evidences obtained in this work. However, the processes of the major interest are all in the cross sections range of microbarns. Therefore, such studies make sense only if the statistical limitations of the existing data are overcome by about an order of magnitude, which is a real challenge for available experimental techniques and facilities.